# An Information-Theoretic Approach to Identifying Formulaic Clusters in Textual Data

### v1

@@ 2.2 Clustering Approach
$s_{i}\in[0,1],\quad s_{i}=0\text{ for full membership in }\mathcal{C}_{1},\quad s_{i}=1\text{ for full membership in }\mathcal{C}_{2}.$ (5)
@@
$\begin{cases}x_{i}\in\mathcal{C}_{1},&\text{if }s_{i}=0,\\ x_{i}\in\mathcal{C}_{2},&\text{if }s_{i}=1,\\ x_{i}\in\mathcal{C}_{1}\text{ and }\mathcal{C}_{2}\text{ with weight }(1-s_{i},s_{i}),&\text{if }0<s_{i}<1.\end{cases}$
@@ 2.3 Clustering Formalism for Discrete Categorical Data
Specifically, the soft probability of the $j$th word appearing in a given cluster is estimated as:
@@
where $\chi$ is the set of unique categorical features satisfying $w\in\chi$, and $p(w|\vec{s})$ represents the probability of category $w$ given a weight sequence $\vec{s}$. The variable $x_{i}$ denotes the feature vector of the $i$th sample. Under the Bernoulli assumption, $x_{ij}=1$ if the $j$th feature is present in the sample and $x_{ij}=0$ otherwise. In the cumulative case, $x_{ij}$ instead represents the number of times the $j$th feature appears in the sample. This formulation allows us to compute empirical word distributions within each cluster while incorporating soft clustering assignments as weighting factors.
@@
Here, $p(w_{j}|\vec{s})$ and $p(w_{j}|1-\vec{s})$ represent the soft Bernoulli probabilities of the $j$th feature being present in each cluster. These probabilities are derived directly from the empirical word distributions calculated in Eq. (7).
@@
While this formulation provides a principled probabilistic basis for clustering, it does not explicitly account for competing structural signals that may emerge from different generative processes within the data. However, assessing the presence and dominance of such alternative signals is beyond the scope of this work, as our primary focus is on the probabilistic clustering framework itself. That said, the relative influence of entropy-based clustering versus clustering driven by feature distributions is inherently dependent on dataset-specific parameters, such as the baseline activation probability of features and the presence of structured formulaic dimensions. A heuristic analysis of this interplay is provided in Appendix C, where we illustrate how these factors affect clustering outcomes and discuss the need for future work to develop a predictive framework for determining when each signal will dominate.
@@ 2.4 Clustering Formalism for Continuous Multivariate Gaussian Data
where $D_{M}(x_{i^{\prime}}|\vec{s})=(x_{i^{\prime}}-\mu_{w})^{\top}\Sigma_{w}^{-1}(x_{i^{\prime}}-\mu_{w})$ is the Mahalanobis distance given the weighting by $\vec{s}$ (see Appendix A). This equation directly relates a sample’s entropy contribution to its likelihood under the dataset’s distribution.
@@
This score combines the standard deviation, the mean, and the inverse of the sample weights’ norm to quantify cluster coherence by measuring variability and predictability. The optimization minimizes variance while maintaining sufficient samples, distinguishing formulaic clusters from non-formulaic ones.
Here, several differences from cross-entropy-based clustering (see §2.3) are noteworthy: (1) Gaussian-based cross-entropy (e.g., Tabor and Spurek, 2014), relies on full covariance estimation, which becomes unstable when the sample size is small relative to the feature dimensionality, leading to singular covariance matrices. Our approach mitigates this issue by refining the score function to enhance the separability of the formulaic cluster. By focusing on isolating this cluster, we effectively reduce reliance on cross-cluster covariance terms, which are typically the most ill-conditioned. This ensures that covariance estimation is primarily constrained to a more compact and homogeneous subset of the data, where the conditioning of the covariance matrix remains more stable. (2) In Gaussian data, formulaic clusters have low self-information due to high predictability. Low entropy in these clusters corresponds to high likelihood, yielding lower negative log-likelihood. In contrast, in discrete categorical data, formulaic clusters are associated with higher self-information due to sparse but predictable occurrences of categories. Thus, Gaussian formulaic clusters exhibit low self-information, while discrete ones exhibit high self-information.
@@ 2.5 Optimization Scheme
The algorithm optimizes weights $\vec{s}\in\mathbb{R}^{n}$ to isolate the auto-entropy distribution of the structured or formulaic layer within the data. The optimization scheme can be summarized as follows:
@@
Below, we define simplified score functions for the Gaussian and Discrete One-Hot cases.
@@ 3.1 Clustering Benchmarking on Categorical Data
However, it is important to acknowledge the inherent difficulty of modeling textual data. Texts are complex, with nuanced structures and dependencies that cannot be fully captured by simplistic binary features or synthetic generation processes. Despite these limitations, the purpose of this benchmark is not to replicate textual data exactly but to demonstrate the ability of our algorithm to solve a difficult nonlinear problem—optimizing the distinction between sparse, discrete datasets that differ in entropy. This serves as a proof-of-concept for the algorithm’s robustness in handling challenging clustering problems.
Here, since insofar as we are aware – no clustering algorithm exists that is tailored to this specific dataset – we consider three nominal clustering algorithms that represent different clustering paradigms: (1) GMM, which assumes data is generated from a mixture of Gaussian distributions, making it suitable for probabilistic clustering in continuous spaces, (2) $k$-means (Hartigan et al., 1979), which partitions data into clusters based on minimizing intra-cluster variance, making it effective for Euclidean distance-based separation, and (3) DBSCAN (Ester et al., 1996), which identifies clusters as dense regions of points separated by low-density areas, making it robust to arbitrary-shaped clusters but sensitive to parameter tuning. The results of this experiment, shown in Fig. 1, demonstrate the superior performance of our algorithm in resolving formulaic clusters in sparse categorical encoded datasets whose sample-to-dimension ratio is small.
@@
As shown in Fig. 1, our method consistently outperforms traditional clustering approaches in the sparse low sample-size-to-dimensions ratio case by effectively distinguishing structured formulaic patterns even in highly sparse discrete encoded data.
3.2 Clustering Benchmarking on Multivariate Gaussian Data
We perform a series of classification experiments to demonstrate the classification power of our algorithm compared to well-established classification routines designed for multivariate Gaussian data. These methods are: (1) Differential-Entropy (Davis and Dhillon, 2006), (2) Gaussian Mixture Expectation-Maximization [^1] https://scikit-learn.org/1.5/modules/mixture.html (GMM), (3) Regularized Expectation-Maximization (Reg. EM) (Houdouin et al., 2023), and (4) Cross-Entropy-Clustering (CEC) (Tabor and Spurek, 2014). Additionally, we test the resolving power of the $L_{2}$ norm distribution, incorporating it into a similar optimization scheme as described for the auto-entropy distribution due to its relevance in high-dimensional settings where the concentration of measure ensures the norm becomes a distinguishing feature (but is independent of the number of samples).
To introduce noise, we modify several eigenvalues of $\Sigma_{1}$ to match those of $\Sigma_{2}$. The noise level is defined as the relative proportion of eigenvalues in $\Sigma_{1}$ replaced with corresponding eigenvalues of $\Sigma_{2}$. In Fig. 2, we present the results of these experiments, showcasing the classification performance of each method under varying sample sizes, numbers of dimensions, and noise levels. The classification accuracy is measured using the Matthews Correlation Coefficient (MCC), given by
@@ 3.2 Clustering Benchmarking on Multivariate Gaussian Data
We find that our classification approach benefits both from the sample size and the number of dimensions and exhibits a considerably weaker sensitivity to a small sample-size-to-dimensions ratio than other classification routines that leverage the multivariate Gaussian assumption. This is because estimating covariance matrices in high-dimensional, small-sample settings is ill-conditioned, leading to unstable classification. Our method circumvents this issue by leveraging auto-entropy distributions rather than direct covariance estimation. Specifically, as shown in Appendix A, self-information-based classification avoids reliance on explicit covariance inversion by computing the soft auto-entropy of individual samples with respect to the weighted dataset entropy.
This formulation improves numerical stability and mitigates singularity issues that arise when the sample size is small relative to the number of dimensions, as demonstrated in Fig. 2.
@@ 4 Application to Textual Data: The Priestly Source(s) in the Pentateuch
The composition of the Pentateuch, or Torah, has been a longstanding subject of critical scholarship, with various theories proposed to account for its linguistic diversity, stylistic inconsistencies, narrative discrepancies, and theological tensions. One of the most widely discussed models is the Documentary Hypothesis, which suggests that the Torah is a composite text formed through the combination of multiple sources, each reflecting different historical, ideological, or scribal traditions (Wellhausen, 1885; Kuenen and Wicksteed, 1886; Holzinger, 1893). Among the sources reconstructed by this framework are the Yahwist (J), the Elohist (E), the Deuteronomist (D), and the Priestly (P) source. While the precise delineation of these sources remains debated, with alternative models proposing alternative groupings or rejecting the hypothesis altogether (Albertz, 2018), the existence of the P material is still widely acknowledged, owing to its distinctive style and content and relative internal consistency (e.g., Kuenen and Wicksteed, 1886; Holzinger, 1893; Haran, 1981; Albertz, 2018). The texts typically classified as P exhibit certain recurring characteristics, including highly structured formulations, genealogical lists, legal material, and a focus on ritual and cultic concerns (e.g., Vink, 1969; Albertz, 2018). Scholars have further proposed that P itself is not monolithic but contains sub-strata, most notably distinguishing between the preceding (or succeeding) Priestly Source (P) and a later Holiness School (H) (e.g., Ross, 1997), which is thought to emphasize purity laws and ethical concerns (e.g., Klostermann, 1907; Knohl, 2007). However, the extent to which these strata represent distinct compositional layers rather than later redactional activity is an ongoing debate, and no consensus exists regarding their precise historical development. Given these complexities, our study does not assume the validity of any particular compositional model but instead seeks to evaluate whether quantitative, entropy-based methods can identify distinct stylistic patterns in the Pentateuch better than previous computerized characterization attempts (Dershowitz et al., 2015; Yoffe et al., 2023; Faigenbaum-Golovin et al., 2024). If such patterns emerge, this may lend support to certain divisions proposed in source-critical scholarship. If they do not, it may suggest that the distinction between the various traditions - particularly, between P and non-P, or P and H – is more fluid than often assumed. We selected P as the main test case for our clustering algorithm due to its highly structured and repetitive nature, which contrasts with the more vivid and colorful style of other Pentateuchal traditions. Because our approach leverages entropy-based clustering to detect formulaic patterns, it provides an opportunity to evaluate whether the hypothesized Priestly texts exhibit distinct formulaic stylistic tendencies that set them apart from surrounding material. Additionally, if a similar distinction between P and H exists (as proposed by Knohl (2007)), our method may provide an independent and complementary means of quantifying and assessing it. Notably, there is broad literary consensus that P constitutes a stylistically distinct body of text, often considered a literary outlier due to its rigid structure, formulaic expressions, and characteristic lexical choices (e.g., Smith, 1996; Bühler et al., 2024). Here, we consider three hypothesized partitions of three books in the Pentateuch for evaluation:
@@
  • Genesis: We test the distinction between genealogical lists and the surrounding narrative material. The great majority of the lists are attributed to P due to their molded expression (e.g., the Toledot formula) and emphasis on lineage continuity (e.g., Long et al., 1984; Plum, 1989). However, a few texts belong to non-P. Hence, the focus in this case is on matters of genre and form in the light of entropy rather than source division. By isolating these passages, we assess whether their auto-entropy properties diverge significantly from the surrounding material. In the second stage, the possibility of internal division between P and non-P within the genealogical lists will be examined/demands further inquiry.
@@
  • Exodus: We analyze the division between Priestly (P) and non-Priestly (non-P) material. Priestly passages in Exodus largely consist of legal and cultic instructions, particularly concerning the construction of the Tabernacle, whereas non-Priestly sections retain a more narrative-driven style (Römer, 2009; Schmid, 2015). This partition has been previously explored in our work (Yoffe et al., 2023, 2024; Bühler et al., 2024), making it a useful benchmark for gauging the performance of our method against previous results.
@@
  • Leviticus: We investigate whether a distinction can be quantitatively detected be- tween the Priestly (P) and Holiness (H) materials. While both exhibit high degrees of formulaicity, the Holiness texts are often argued to introduce distinctive stylistic and theological elements (Knohl, 2007). If our algorithm can recover a meaningful distinction, this may provide additional empirical support for the hypothesized P/H division.

### v2

@@ 2.2 Clustering Approach
[…]
@@
[…]
@@
Layman explanation:
Rather than forcing each text unit of text into one cluster or the other, our method allows for partial or uncertain membership between the two. This flexibility is important in cases where text units may contain a mix of formulaic and non-formulaic features. During optimization, each segment is evaluated based on how structurally predictable it is. Highly predictable text units lean toward the formulaic cluster, while less predictable ones lean toward the non-formulaic cluster. This “soft” classification makes the method more stable and better suited for analyzing noisy or stylistically mixed texts.
@@ 2.3 Clustering Formalism for Discrete Categorical Data
Layman explanation:
The goal is to ensure that the feature distributions within each cluster align well with the actual data. Cross-entropy serves as a standard measure of divergence between observed and modeled probabilities. Minimizing it improves the model’s ability to describe how features are distributed in the data.
@@
Layman explanation:
The Bernoulli model treats each feature as either present or absent in a given segment of text. This is well-suited to binary or count-based feature representations. We estimate the frequency of each feature within a cluster by aggregating across samples, weighted by how strongly they belong to the cluster.
Specifically, the soft probability of the $j$-th word appearing in a given cluster is estimated as:
@@
where $\chi$ is the set of unique categorical features satisfying $w\in\chi$, and $p(w|\vec{s})$ represents the probability of category $w$ given a weight sequence $\vec{s}$. The variable $x_{i}$ denotes the feature vector of the $i$-th sample. Under the Bernoulli assumption, $x_{ij}=1$ if the $j$-th feature is present in the sample and $x_{ij}=0$ otherwise. In the cumulative case, $x_{ij}$ instead represents the number of times the $j$-th feature appears in the sample. This formulation allows us to compute empirical word distributions within each cluster while incorporating soft clustering assignments as weighting factors.
Layman explanation:
This step produces a probability distribution over features for each cluster. These distributions capture what makes each cluster statistically distinct, based on how often specific features appear in text segments that are likely to belong to that cluster.
@@
Here, $p(w_{j}|\vec{s})$ and $p(w_{j}|1-\vec{s})$ represent the soft Bernoulli probabilities of the $j$-th feature being present in each cluster. These probabilities are derived directly from the empirical word distributions calculated in Eq. (7).
Layman explanation:
These expressions compute how likely a full set of features is for a given sample under each cluster model. If a text segment contains many features that are common in one cluster but rare in the other, the value will reflect that asymmetry.
@@
Layman explanation:
The score function evaluates how well the current cluster assignments explain the data. The algorithm updates these assignments to minimize the score, which corresponds to improving the model’s alignment with the observed feature patterns.
@@
Layman explanation:
When features can appear multiple times in a sample, such as when counting words as many times as they appear in a text unit, rather than whether the word appears at all or not, as in the binary case, the Binomial model accounts for these counts. This extension preserves the logic of the model while improving its fit to data with elements that are counted repeatedly.
@@
Layman explanation:
This final expression combines all the elements described above into a unified scoring objective. The clustering process then proceeds by adjusting the sample weights to find the configuration that most efficiently captures the structural differences in the data.
While this formulation provides a principled probabilistic basis for clustering, it does not explicitly account for competing structural signals that may emerge from different generative processes within the data. However, assessing the presence and dominance of such alternative signals is beyond the scope of this work, as our primary focus is on the probabilistic clustering framework itself. That said, the relative influence of entropy-based clustering versus clustering driven by feature distributions is inherently dependent on dataset-specific parameters, such as the baseline activation probability of features and the presence of structured formulaic dimensions. A heuristic analysis of this interplay is provided in C, where we illustrate how these factors affect clustering outcomes and discuss the need for future work to develop a predictive framework for determining when each signal will dominate.
@@ 2.4 Clustering Formalism for Continuous Multivariate Gaussian Data
Layman explanation:
In the discrete case, each text segment is represented by binary indicators, capturing the presence or absence of individual features (e.g., words or $n$-grams) in a text unit. However, this representation does not extend to settings where features are continuous-valued, such as neural embeddings or normalized frequency vectors. In these cases, one must model how the values themselves vary across samples rather than simply whether a feature occurs. The multivariate Gaussian distribution provides a natural and mathematically grounded way to describe such continuous variation. Within this framework, entropy reflects how concentrated or dispersed the distribution is, offering a principled measure of structural regularity in continuous feature space.
@@
where $D_{M}(x_{i^{\prime}}|\vec{s})=(x_{i^{\prime}}-\mu_{w})^{\top}\Sigma_{w}^{-1}(x_{i^{\prime}}-\mu_{w})$ is the Mahalanobis distance given the weighting by $\vec{s}$ (see A).
Layman explanation:
A sample’s self-information increases the further it lies from the mean of the data, especially if it does so along directions where the data varies less. This distance, measured by the Mahalanobis metric, provides a principled way of quantifying how typical or atypical a given point is relative to the distribution of other points.
This equation directly relates a sample’s entropy contribution to its likelihood under the dataset’s distribution.
@@
Layman explanation:
The score function favors clusters that are internally consistent (low variability), strongly predictable (low average self-information), and include enough samples to be meaningful. These three criteria together help isolate regions of the data that exhibit formulaic regularity.
This score combines the standard deviation, the mean, and the inverse of the sample weights’ norm to quantify cluster coherence by measuring variability and predictability. The optimization minimizes variance while maintaining sufficient samples, distinguishing the formulaic cluster. Here, several differences from cross-entropy-based clustering (see §2.3) are noteworthy: (1) Gaussian-based cross-entropy (tabor2014cross, 60, e.g.,), relies on full covariance estimation, which becomes unstable when the sample size is small relative to the feature dimensionality, leading to singular covariance matrices. Our approach mitigates this issue by refining the score function to enhance the separability of the formulaic cluster. By focusing on isolating this cluster, we effectively reduce reliance on cross-cluster covariance terms, which are typically the most ill-conditioned. This ensures that covariance estimation is primarily constrained to a more compact and homogeneous subset of the data, where the conditioning of the covariance matrix remains more stable. (2) In Gaussian data, formulaic clusters have low self-information due to high predictability. Low entropy in these clusters corresponds to high likelihood, yielding lower negative log-likelihood. In contrast, in discrete categorical data, formulaic clusters are associated with higher self-information due to sparse but predictable occurrences of categories. Thus, Gaussian formulaic clusters exhibit low self-information, while discrete ones exhibit high self-information.
@@ 2.5 Optimization Scheme
The algorithm optimizes weights $\vec{s}\in\mathbb{R}^{n}$ to isolate the self-information distribution of the structured or formulaic layer within the data. The optimization scheme is summarized in Algorithm 1. While our method is flexible and conceptually transparent, it is also more computationally intensive than conventional iterative clustering algorithms such as $k$-means or Expectation-Maximization (EM). Those methods benefit from closed-form updates for centroids or covariance matrices, leading to fast convergence in low-dimensional, well-behaved data. In contrast, our approach involves nonlinear optimization over a continuous weight vector. This added cost reflects the complexity of modeling self-information and structural predictability, rather than geometric proximity or global variance. In scenarios where sample sizes are large and feature dimensionality is low, iterative methods may indeed be more efficient and comparably effective. However, this is rarely the case in historical or literary corpora, which tend to exhibit sparse, high-dimensional feature spaces and relatively small sample counts. It is precisely these conditions that motivated the design of an information-theoretic clustering approach tailored to such data.
@@ 3.1 Clustering Benchmarking on Categorical Data
However, it is important to acknowledge the inherent difficulty of modeling textual data. Texts are complex, with nuanced structures and dependencies that cannot be fully captured by simplistic binary features or synthetic generation processes. Despite these limitations, the purpose of this benchmark is not to replicate textual data exactly but to demonstrate the ability of our algorithm to solve a difficult nonlinear problem—optimizing the distinction between sparse, discrete datasets that differ in entropy. This serves as proof of the algorithm’s robustness in handling challenging clustering problems.
We consider three nominal clustering algorithms that represent different clustering paradigms: (1) GMM, which assumes data is generated from a mixture of Gaussian distributions, making it suitable for probabilistic clustering in continuous spaces, (2) $k$-means hartigan1979k (23), which partitions data into clusters based on minimizing intra-cluster variance, making it effective for Euclidean distance-based separation, and (3) DBSCAN ester1996density (16), which identifies clusters as dense regions of points separated by low-density areas, making it robust to arbitrary-shaped clusters but sensitive to parameter tuning. The results of this experiment, shown in Figure 1, demonstrate the superior performance of our algorithm in resolving formulaic clusters in sparse categorical encoded datasets whose sample-to-dimension ratio is small.
@@
The classification accuracy is measured using the Matthews Correlation Coefficient (MCC), given by
@@
As shown in Figure 1, our method consistently outperforms traditional clustering approaches in the sparse low sample-size-to-dimensions ratio case by effectively distinguishing structured formulaic patterns even in highly sparse discrete encoded data.
3.2 Clustering Benchmarking on Multivariate Gaussian Data
We perform a series of classification experiments to demonstrate the classification power of our algorithm compared to well-established classification routines designed for multivariate Gaussian data. These methods are: (1) Differential-Entropy davis2006differential (12), (2) Gaussian Mixture Expectation-Maximization [^1] https://scikit-learn.org/1.5/modules/mixture.html (GMM), (3) Regularized Expectation-Maximization (Reg. EM) houdouin2023regularized (25), and (4) Cross-Entropy-Clustering (CEC) tabor2014cross (60). Additionally, we test the resolving power of the $L_{2}$ norm distribution, incorporating it into a similar optimization scheme as described for the self-information distribution due to its relevance in high-dimensional settings where the concentration of measure ensures the norm becomes a distinguishing feature (but is independent of the number of samples).
To introduce noise, we modify several eigenvalues of $\Sigma_{1}$ to match those of $\Sigma_{2}$. The noise level is defined as the relative proportion of eigenvalues in $\Sigma_{1}$ replaced with corresponding eigenvalues of $\Sigma_{2}$. In Figure 2, we present the results of these experiments, showcasing the classification performance of each method under varying sample sizes, numbers of dimensions, and noise levels.
@@ 3.2 Clustering Benchmarking on Multivariate Gaussian Data
We find that our classification approach benefits both from the sample size and the number of dimensions and exhibits a considerably weaker sensitivity to a small sample-size-to-dimensions ratio than other classification routines that leverage the multivariate Gaussian assumption. This is because estimating covariance matrices in high-dimensional, small-sample settings is ill-conditioned, leading to unstable classification. Our method circumvents this issue by leveraging auto-entropy distributions rather than direct covariance estimation. Specifically, as shown in A, self-information-based classification avoids reliance on explicit covariance inversion by computing the soft auto-entropy of individual samples with respect to the weighted dataset entropy.
This formulation improves numerical stability and mitigates singularity issues that arise when the sample size is small relative to the number of dimensions, as illustrated in Figure 2.
@@ 4 Application to Textual Data: The Priestly Source(s) in the Pentateuch
The composition of the Pentateuch, or Torah, has been a longstanding subject of critical scholarship, with various theories proposed to account for its linguistic diversity, stylistic inconsistencies, narrative discrepancies, and theological tensions. One of the most widely discussed models is the Documentary Hypothesis, which suggests that the Torah is a composite text formed through the combination of multiple sources, each reflecting different historical, ideological, or scribal traditions wellhausen1885prolegomena (64, 34, 24). Among the sources reconstructed by this framework are the Yahwist (J), the Elohist (E), the Deuteronomist (D), and the Priestly (P) source. While the precise delineation of these sources remains debated, with alternative models proposing alternative groupings or rejecting the hypothesis altogether albertz2018recent (1), the existence of the P material is still widely acknowledged, owing to its distinctive style and content and relative internal consistency (kuenen1886historico, 34, 24, 22, 1, e.g.,). The texts typically classified as P exhibit certain recurring characteristics, including highly structured formulations, genealogical lists, legal material, and a focus on ritual and cultic concerns (vink1969date, 61, 1, e.g.,). Scholars have further proposed that P itself is not monolithic but contains sub-strata, most notably distinguishing between the preceding (or succeeding) Priestly Source (P) and a later Holiness School (H) (ross1997composition, 51, e.g.,), which is thought to emphasize purity laws and ethical concerns (klostermann1907pentateuch, 32, 33, e.g.,). However, the extent to which these strata represent distinct compositional layers rather than later redactional activity is an ongoing debate, and no consensus exists regarding their precise historical development. Given these complexities, our study does not assume the validity of any particular compositional model but instead seeks to evaluate whether quantitative, entropy-based methods can identify distinct stylistic patterns in the Pentateuch better than previous computerized characterization attempts dershowitz2015computerized (13, 68, 17). If such patterns emerge, this may lend support to certain divisions proposed in source-critical scholarship. If they do not, it may suggest that the distinction between the various traditions, particularly between P and non-P, or P and H, is more fluid than often assumed. We selected P as the main test case for our clustering algorithm due to its highly structured and repetitive nature, which contrasts with the more vivid and colorful style of other Pentateuchal traditions. Because our approach leverages entropy-based clustering to detect formulaic patterns, it provides an opportunity to evaluate whether the hypothesized Priestly texts exhibit distinct formulaic stylistic tendencies that set them apart from surrounding material. Additionally, if a similar distinction between P and H exists (as proposed by knohl2007sanctuary (33)), our method may provide an independent and complementary means of quantifying and assessing it. Notably, there is a broad literary consensus that P constitutes a stylistically distinct body of text, often considered a literary outlier due to its rigid structure, formulaic expressions, and characteristic lexical choices (smith1996literary, 57, 6, e.g.,). Here, we consider three hypothesized partitions of the three books in the Pentateuch for evaluation:
@@
  • Genesis: We test the distinction between genealogical lists and the surrounding narrative material. The great majority of the lists are attributed to P due to their molded expression (e.g., the Toledot formula) and emphasis on lineage continuity (long1984toledotformel, 37, 43, e.g.,). However, a few texts belong to non-P. Hence, the focus in this case is on matters of genre and form in the light of entropy rather than source division. By isolating these passages, we assess whether their self-information properties diverge significantly from the surrounding material. In the second stage, the possibility of internal division between P and non-P within the genealogical lists will be examined/demands further inquiry.
@@
  • Exodus: We analyze the division between Priestly (P) and non-Priestly (non-P) material. Priestly passages in Exodus largely consist of legal and cultic instructions, particularly concerning the construction of the Tabernacle, whereas non-Priestly sections retain a more narrative-driven style romer2009exodus (50, 53). This partition has been previously explored yoffe2023statistical (68, 6), making it a useful benchmark for gauging the performance of our method against previous results.
@@
  • Leviticus: We investigate whether a distinction can be quantitatively detected between the Priestly (P) and Holiness (H) materials. While both exhibit high degrees of formulaicity, the Holiness texts are often argued to introduce distinctive stylistic and theological elements knohl2007sanctuary (33). If our algorithm can recover a meaningful distinction, this may provide additional empirical support for the hypothesized P/H division.
Each of these test cases represents a distinct type of contrast in structural predictability. In Genesis, the distinction between genealogical lists and surrounding narrative material constitutes the most pronounced formulaicity contrast, as the genealogies are highly repetitive and rigid in structure. In Exodus, the contrast between Priestly and non-Priestly material also exhibits a clear difference, particularly in sections related to cultic legislation and Tabernacle construction, which display strongly regularized patterns yoffe2023statistical (68, 6). The P/H distinction in Leviticus had not previously been analyzed using computerized methods. Its inclusion here allows us to assess whether subtle stylistic differences, such as those associated with the Holiness corpus, manifest in statistically measurable differences in self-information. Together, these examples are chosen to gauge the method’s ability to quantify varying degrees of formulaicity across multiple strata of the biblical text.